Question 15 Previous Next> Assume a companys Income Statement for Year 12 is as follows Year 12 in 000s Income Statement Data Net Revenues from Footwear Sales Cost of Pairs Sold Warehouse Expenses Marketing Expenses Administrative Expenses Operating Profit (Loss) Interest Income (Expense) Pre-tax Profit (Loss) Income Taxes Net Profit (Loss $530,000 340,000 40,000 80,000 15,000 55,000 (10,000) 45,000 13,500 S 31,500 Based on the above income statement data and assuming the company has 20 million shares of common stock outstanding, the companys operating profit margin and EPS were O 5.94% and $1.55 O 6.67% and $1.84 10.4% and $2.75 8.49% and $2.25 O 10.38% and $1.55 Copying, redistributing, or website posting is expressly prohibited and constitutes copyright violation. Copyright 2018 by Glo-Bus Software, Inc. < Previous Next>
Answer #1

Answer 15:

  • option 5: 10.38% and $1.55

Explanation:

  • operating profit = 55000
  • Revenue = 530000
  • Operating margin = operating profit / revenue = 55000/530000 = 0.10377= 10.38%

  • net profit = 31500 (000)
  • share = 20 million
  • EPS = net profit / shares = $1.55

Note: Chegg mandates answering of first full question, please repost the other question to get them answered by an expert.